Grade 12Mechanics

A metallic sphere weighs 35g in air and 28.5g in water. Find its relative density and also findits weight in a liquid of relative density 0.9

nidhi
8 Years agoGrade 12

2 Answers

Turbashu Saha
6 Years ago
Given wt. of sphere in air = 35 g
Now, wt. of water of volume equal to that of sphere = apparent loss in weight of sphere 
 
I.e. weight of water of volume equal to that of sphere = 35-28.5 =6.5 g
 
Therefore relative density = 35/6.5 = 5.4 
 
Now density of water is 1g/cc
 
Therefore,  density of solid sphere = 5.4×1 = 5.4 g/cc
 
Therefore its volume = 35/5.4 =6.5 g/cc
 
Now , apparent wt in 2nd liquid = Actual wt- loss in wt in liquid 
 
= (35- 6.5×0.9) 
=29.15 g
Vikas TU
6 Years ago
Dear student 
 density of solid sphere = 5.4×1 = 5.4 g/cc 
Density of water = 1 g /cc 
So relative density = 5.4 / 1 = 5.4 
Hope this will help